OpenStudy (anonymous):

The sum of two numbers is 53 . The smaller number is 7 less than the larger number. What are the numbers?

OpenStudy (anonymous):

@NietzscheAlex

OpenStudy (anonymous):

let larger no.=x hence smaller no.=x-7 ATQ X+X-7=53 2X=53-7 X=23

OpenStudy (anonymous):

oh okay that's the formula and thank you @lall

OpenStudy (anonymous):

SORRY ther is mistake 2x=60 x=30

OpenStudy (anonymous):

largest no. =30 smallest no. is 30-7=23

OpenStudy (anonymous):

checking the ans. 30+23=53

OpenStudy (anonymous):

@lall so that's how I have to set it up

OpenStudy (anonymous):

yes see the statements and form the eqn.'s

OpenStudy (anonymous):

oh okay thank you and well can you help me with this one its tricky for me @lall